The terminal side of theta, an angle in standard position, is in Quadrant II and lies on the line y=-8x. What are the six trig functions for this line?
Answer by stanbon(75887)

The terminal side of theta, an angle in standard position, is in Quadrant II and lies on the line y=-8x. What are the six trig functions for this line?
-----------------------
Point on the line: (-1,8)
----
You have x = -1 and y = 8
Find "r":
r^2 = (-1)^2+8^2
r^2 = 65
r = sqrt(65)
==================
sin = y/r = 8/sqrt(65)
cos = -1/sqrt(65)
tan = y/x = -8
----
csc = r/y = sqrt(65)/8
sec = r/x = -sqrt(65)
ctn = x/y = -1/8
================
Cheers,
Stan H.
================